# Paul-Erik Haacker, Remco I. Leine, Renu Chaudhary, Kai Diethelm, André Schmidt, Safoura Hashemishahraki

### v1

@@ 6 Conclusion and Open Questions
    In the case of LTP-FODEs, how can one construct a monodromy operator and is it possible to find Floquet multipliers? Since the history of trajectories is a part of FODEs, we conjecture that a monodromy operator $\Phi_{T}:X(-\infty,0)\to X(-\infty,T)$ is required for the study of LTP systems analogous to the case of functional differential equations [22]. The spectrum of this operator will map the stability properties of the respective system. Consequently, the study of fundamental solution operators for LTP-FODEs is a neccessary first step.
@@ Acknowledgements
Assume w.l.o.g. $t_{0}=0$. We first show that that $\mathcal{F}x_{0}\in C[0,\infty)$. In other words, we must demonstrate that, for all $t\in[0,\infty)$, $\mathcal{F}x(t+\varepsilon)\to\mathcal{F}x(t)$ as $\varepsilon\to 0$.
@@
$\mathcal{F}x(t+\varepsilon)-\mathcal{F}x(t)=\frac{1}{\Gamma(1-\alpha)}(B_{1}+B_{2})$ (85)
@@
$B_{1}$ $=\int_{-\eta}^{0}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]x^{\prime}(\tau)\,\mathrm{d}\tau$ (86)
@@
$B_{2}$ $=\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]x^{\prime}(\tau)\,\mathrm{d}\tau$ (87)
where $\eta>0$ is the fixed number from the property that $x^{\prime}\in L^{\infty}[-\eta,0]$. Recalling that $x^{\prime}$ is bounded on $\eta,0]$, we see that
$|B_{1}|$ $\leq\sup_{\tau\in[-\eta,0]}|x^{\prime}(\tau)|\cdot\int_{-\eta}^{0}\left|(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right|\,\mathrm{d}\tau$ (88)
$=\sup_{\tau\in[-\eta,0]}|x^{\prime}(\tau)|\frac{1}{1-\alpha}\left[(t+\eta)^{1-\alpha}-(t+\varepsilon+\eta)^{1-\alpha}-t^{1-\alpha}+(t+\varepsilon)^{1-\alpha}\right],$ (89)
where we used $(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}<0$ and therefore $B_{1}\to 0$ for $\varepsilon\to 0$. Moreover, an integration by parts yields that
@@
$B_{21}$ $=\left.\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]^{\vphantom{b}}\cdot x(\tau)\right|_{\tau=-\infty}^{\tau=-\eta}$ (91)
@@
$B_{22}$ $=-\alpha\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right]x(\tau)\,\mathrm{d}\tau.$
Clearly, since $x$ is bounded and the term in square brackets in (91) tends to zero as $\tau\to$-$\infty$, we find that
$B_{21}=\left[(t+\varepsilon+\eta)^{-\alpha}-(t+\eta)^{-\alpha}\right]\cdot x(-\eta)\to 0$ (92)
for $\varepsilon\to 0$ since $t\geq 0$ and $\eta>0$. Finally, once again using the boundedness of $x$ and applying the mean value theorem of differential calculus to the function $(\cdot)^{-1-\alpha}$ on the interval $t-\tau,t+\varepsilon-\tau]$, we see that
$|B_{22}|$ $\leq\alpha\sup_{\tau\leq-\eta}|x(\tau)|\int_{-\infty}^{-\eta}\left|(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right|\,\mathrm{d}\tau$ (93)
$=\alpha(\alpha+1)\sup_{\tau\leq-\eta}|x(\tau)|\cdot\varepsilon\int_{-\infty}^{-\eta}\left|\xi_{\tau,\varepsilon}\right|^{-2-\alpha}\,\mathrm{d}\tau$ (94)
@@
$|B_{22}|$ $\leq\varepsilon~\alpha(\alpha+1)\sup_{\tau\leq-\eta}|x(\tau)|\cdot\int_{-\infty}^{-\eta}(t-\tau)^{-2-\alpha}\,\mathrm{d}\tau$ (95)
@@
which does hold since $\zeta\in[0,1)$. Thus, equation (100) indeed holds. For the second integral we apply integration by parts,

### v2

@@ 6 Conclusion and Open Questions
    In the case of LTP-FODEs, how can one construct a monodromy operator and is it possible to find Floquet multipliers? Since the history of trajectories is a part of FODEs, we conjecture that a monodromy operator $\Phi_{T}:X(-\infty,0)\to X(-\infty,T)$ is required for the study of LTP systems analogous to the case of functional differential equations [Hale1977]. The spectrum of this operator will map the stability properties of the respective system. Consequently, the study of fundamental solution operators for LTP-FODEs is a neccessary first step.
@@ Acknowledgements
Assume w.l.o.g. $t_{0}=0$. We first show that that $\mathcal{F}x_{0}\in C[0,\infty)$. In other words, we must demonstrate that, for all $t\in[0,\infty)$, $\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t+\varepsilon)\to\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t)$ as $\varepsilon\to 0$.
@@
$\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t+\varepsilon)-\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t)=\frac{1}{\Gamma(1-\alpha)}(B_{1}+B_{2})$ (87)
@@
$B_{1}$ $=\int_{-\eta}^{0}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (88)
@@
$B_{2}$ $=\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (89)
where $\eta>0$ is the fixed number from the property that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}\in L^{\infty}[-\eta,0]$. Recalling that $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$$^{\prime}$ is bounded on $\eta,0]$, we see that
$|B_{1}|$ $\leq\sup_{\tau\in[-\eta,0]}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)|\cdot\int_{-\eta}^{0}\left|(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right|\,\mathrm{d}\tau$ (90)
$=\sup_{\tau\in[-\eta,0]}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)|\frac{1}{1-\alpha}\left[(t+\eta)^{1-\alpha}-(t+\varepsilon+\eta)^{1-\alpha}-t^{1-\alpha}+(t+\varepsilon)^{1-\alpha}\right],$ (91)
where we used $(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}<0$ and therefore $B_{1}\to 0$ for $\varepsilon\to 0$. Moreover, an integration by parts, justified by the absolute continuity of $x_{0}$, yields that
@@
$B_{21}$ $=\left.\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]^{\vphantom{b}}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)\right|_{\tau=-\infty}^{\tau=-\eta}$ (93)
@@
$B_{22}$ $=-\alpha\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)\,\mathrm{d}\tau.$
Clearly, since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$} is bounded and the term in square brackets in (93) tends to zero as $\tau\to$-$\infty$, we find that
$B_{21}=\left[(t+\varepsilon+\eta)^{-\alpha}-(t+\eta)^{-\alpha}\right]\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(-\eta)\to 0$ (94)
for $\varepsilon\to 0$ since $t\geq 0$ and $\eta>0$. Finally, once again using the boundedness of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$} and applying the mean value theorem of differential calculus to the function $(\cdot)^{-1-\alpha}$ on the interval $t-\tau,t+\varepsilon-\tau]$, we see that
$|B_{22}|$ $\leq\alpha\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\int_{-\infty}^{-\eta}\left|(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right|\,\mathrm{d}\tau$ (95)
$=\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\cdot\varepsilon\int_{-\infty}^{-\eta}\left|\xi_{\tau,\varepsilon}\right|^{-2-\alpha}\,\mathrm{d}\tau$ (96)
@@
$|B_{22}|$ $\leq\varepsilon~\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\cdot\int_{-\infty}^{-\eta}(t-\tau)^{-2-\alpha}\,\mathrm{d}\tau$ (97)
@@
which does hold since $\zeta\in[0,1)$. Thus, equation (102) indeed holds. For the second integral we apply integration by parts, justified by the absolute continuity of $x_{0}$,